OpenStudy (anonymous):

simplify 5/x^2+1/x^2+x=???

OpenStudy (anonymous):

I could not have answer for it \[\frac{ 5 }{ x^2 }+\frac{ 1 }{ x^2+x }=?\]

OpenStudy (baru):

take common denominator \[\frac{ 5(x^2+x) + x^2 }{ x^2(x^2+x) }\]

OpenStudy (baru):

\[\frac{ 5x^2+5x + x^2 }{ (x^4+x^3) }\]

OpenStudy (anonymous):

yes i did that one \[\frac{ x(6x+5) }{ x^2(x^2+x) }\]

OpenStudy (anonymous):

but when i Puget the said wrong

OpenStudy (baru):

the 'x' in the numerator and denominator cancel \[\frac{ 6x+5 }{ (x^3+x^2) }\]

OpenStudy (anonymous):

ok i will do it thanks